# GLIMPSE-D: An Exotic Balmer-Jump Object at ? Revisiting Photometric Selection and the Cosmic Abundance of PopĀ III Galaxies

### v1

@@ 5.3 Updated PopĀ III UVLF and CSFRD
FigureĀ 9 presents the updated PopĀ III UVLF. The red filled pentagon denotes the spectroscopic candidate of AMORE6, and the red arrows mark the upper limits in the other magnitude bins. For the spec- candidates, the constraint remains broadly consistent with theoretical expectations from recent simulations (visbal2020; liu2020; sarmento2022; venditti2023; venditti2025). The corresponding number densities in each magnitude bin are summarized in TableĀ 3.
@@ 6 Summary
We performed 30.4Ā hr of on-source NIRSpec/G395M observations. The spectra clearly detect H, H, and [Oāiii], with line ratios of H/H and [Oāiii]/H. These measurements show that GLIMPSE-16043 is not a genuine zero-metallicity PopĀ III system.
@@
By combining the line fluxes with NIRCam photometry, we constrain the underlying continuum and find an extraordinarily strong Balmer jump ( mag) and an extreme rest-frame H EW (āĆ ). These values exceed predictions from all current stellar+nebular and pure nebular photoionization models.
@@
We test a suite of models: (A) BPASS+Cloudy and (B) PopĀ III+Cloudy reproduce some qualitative trends with the strong EW(H) but fail to match the Balmer jump; (C) pure nebular continua cannot simultaneously fit the Balmer jump and UV slope; (D) blackbody+Cloudy models with K enhance [Oāiii] but weaken the Balmer jump; and (E) blackbody+Cloudy models with fixed provide the closest match with a K blackbody and K, yielding a Balmer jump amplitude of mag, still insufficient to reproduce all observed properties. This points toward more exotic origins.
@@
Appendix A Updated photometry of GLIMPSE-16043
@@ Appendix C Tentative Candidates
To further characterize the system, we fix the Ly redshift and flux and perform SED modeling with Bagpipes (carnall2018) for each of the five clumps; their best-fit SEDs are shown in FigureĀ 14. The Bagpipes fits indicate that A370-z6LAE-1c represents a low-mass () stellar cluster hosted by a galaxy, and they confirm that clumpĀ exhibits very weak [Oāiii] emission, in contrast to the strong [Oāiii] seen in most of the other clumps. The weak [Oāiii] is also independently supported by the latest deep F360M imaging in this field (#5890; PIs: C.Ā Withers & A.Ā Muzzin), which shows no enhancement relative to F356W. Conversely, the NIRCam F360M grism data obtained in the same region (PIDĀ 2883; PI: F.Ā Sun) show marginal positive flux at the expected [Oāiii] wavelength inferred from the Ly redshift (integrated S/N 3), suggesting the possibility of non-zero [Oāiii] emission in A370-z6LAE-1c. A key caution for both the imaging and grism measurements of A370-z6LAE-1c is the presence of a nearby bright cluster-member galaxy, which makes the results highly sensitive to the details of local-background subtraction. In fact, an independent reduction and analysis of the F360M grism data yield the possible [Oāiii] feature much weaker with S/N , highlighting the level of practical uncertainty introduced by bright neighboring sources. Taken together, these results reinforce the tentative nature of A370-z6LAE-1c as a PopĀ III candidate, while simultaneously demonstrating how strong gravitational lensing enables us to probe distinct stellar populations inside a galaxy.
@@
A370-z6LAE-2 satisfies all of our updated PopIII selection criteria in both the colorācolor diagrams and the SED-based tests, except for the requirement of (Eq.7). Its measured S/N in F410M is only 1.4, which naturally reflects the faintness of the object. As discussed in Sec.Ā 5.1, this specific F410M requirement is designed to exclude contaminants exhibiting a significant Balmer jump, such as the case encountered for GLIMPSE-16043. Although the current position of A370-z6LAE-2 in FigureĀ 6 lies well away from the region occupied by the exotic Balmer-jump contaminants, the low S/N in F410M permits its inferred location to scatter into that regime. On the other hand, A370-z6LAE-2 is relatively isolated, with no bright neighboring sources, implying that deeper F410M imaging alone, even without the need for spectroscopy, could robustly establish its color and elevate it to a promising PopĀ III candidate.
@@ Appendix D Survey Area
For each cluster, we compute the source-plane area as a function of magnification , binned logarithmically from to . At each pixel inside the defined region, the magnification is obtained from the lens model (assuming a source redshift of ), and the corresponding source-plane contribution is computed as the image-plane pixel area divided by . Pixels are then accumulated into cumulative magnification bins, such that each bin records the total source-plane area with magnification exceeding a given threshold, . We then convert the ā relation into a relation between and the effective limiting magnitudes for each field. The limiting depths are obtained by scaling according to the magnification and by adopting the F200W sensitivities reported in the literature for the lensing fields analyzed in this study. For non-lensed general blank fields, we use the detection limits and survey areas compiled in fujimoto2025. The resulting ādepth relations for all fields are presented in FigureĀ 16. These relations are used directly to compute the magnification-corrected survey volume employed in our updated PopĀ III UVLF and CSFRD analysis (SectionĀ 5.3).

### v3

@@ 4.3 Modeling GLIMPSE-16043
This requires a situation where PopĀ III stars reside in metal-enriched gas during a āself-pollutedā phase, following the first supernovae (e.g., fwang2012; rusta2025).
@@ 5.3 Updated PopĀ III UVLF and CSFRD
FigureĀ 9 presents the updated PopĀ III UVLF. The red filled pentagon denotes the spectroscopic candidate AMORE6, and the red arrows indicate the upper limits in the remaining magnitude bins. Our constraints, including the upper limits, remain broadly consistent with theoretical expectations from recent simulations (visbal2020; liu2020; sarmento2022; venditti2023; venditti2025). Future large lensing cluster surveys with JWST (e.g., VENUS; fujimoto2025_6882) will help disentangle these possible models and provide a deeper understanding of the first episodes of star formation in the early universe. The corresponding number densities in each magnitude bin are summarized in TableĀ 3.
@@
The upper limit is estimated from the Poisson uncertainty at the single-sided confidence level of 84.13% presented in gehrels1986.
@@ 6 Summary
We performed 30.4Ā hr of on-source NIRSpec/G395M observations of GLIMPSE-16043. The spectra clearly detect H, H, and [Oāiii], with line ratios of H/H and [Oāiii]/H. These measurements show that GLIMPSE-16043 is not a genuine zero-metallicity PopĀ III system.
@@
By combining the NIRSpec line fluxes with NIRCam photometry, we constrain the underlying continuum and find an extraordinarily strong Balmer jump ( mag) and an extreme rest-frame H equivalent width (āĆ ). Both values exceed the predictions of standard stellar+nebular photoionization models and explain why GLIMPSE-16043 was initially misidentified as a promising PopĀ III candidate.
@@
We test a suite of physically motivated models to account for the extreme SED of GLIMPSE-16043: (A) BPASS+Cloudy and (B) PopĀ III+Cloudy reproduce some qualitative trends with the strong EW(H) but fail to match the Balmer jump; (C) pure nebular continua cannot simultaneously fit the Balmer jump and UV slope; (D) blackbody+Cloudy models with K enhance [Oāiii] but weaken the Balmer jump; and (E) blackbody+Cloudy models with fixed provide the closest match with a K blackbody and K, yielding a Balmer jump amplitude of mag, still insufficient to reproduce all observed properties. This points toward more exotic origins.
@@
Acknowledgements.
We are grateful to Kohei Inayoshi and Pratika Dayal for helpful discussions for the physical interpretation of GLIMPSE-16043, Adam Muzzin and Fengwu Sun for useful discussions, inputs, and sharing the data to investigate tentative candidates found in the AbellĀ 370 field, and Takahiro Morishita for sharing the latest photometry of AMORE6-B.This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #9223. Support for program #9223 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. SF acknowledges support from the Dunlap Institute, funded through an endowment established by the David Dunlap family and the University of Toronto. AV acknowledges funding from the Cosmic Frontier Center and the University of Texas at Austinās College of Natural Sciences.Appendix A Updated photometry of GLIMPSE-16043
@@ Acknowledgements.
The measurements above are aperture (-diameter) corrected photometry derived from PSF-matched images in the observed frame (i.e., without correcting for magnification).
@@ Appendix C Tentative Candidates
To further characterize the system, we fix the Ly redshift and flux and perform SED modeling with Bagpipes (carnall2018) for each of the five clumps; their best-fit SEDs are shown in FigureĀ 14. The Bagpipes fits indicate that A370-z6LAE-1c represents a low-mass () stellar cluster hosted by a galaxy, and they confirm that clumpĀ exhibits very weak [Oāiii] emission, in contrast to the strong [Oāiii] seen in most of the other clumps. The weak [Oāiii] is also independently supported by the latest deep F360M imaging in this field (#5890; PIs: C.Ā Withers & A.Ā Muzzin), which shows no enhancement relative to F356W. Conversely, the NIRCam F360M grism data obtained in the same region (PIDĀ 2883; PI: F.Ā Sun) show marginal positive flux at the expected [Oāiii] wavelength inferred from the Ly redshift (integrated S/N 3), suggesting the possibility of non-zero [Oāiii] emission in A370-z6LAE-1c. A key caution for both the imaging and grism measurements of A370-z6LAE-1c is the presence of a nearby bright cluster-member galaxy, which makes the results highly sensitive to the details of local-background subtraction. Indeed, an independent reduction and analysis of the F360M grism data yield a weaker possible [Oāiii] feature with S/N , highlighting the level of practical uncertainty introduced by bright neighboring sources. Taken together, these results reinforce the tentative nature of A370-z6LAE-1c as a PopĀ III candidate, while simultaneously demonstrating the natural importance of the spatially-resolved approach and how strong gravitational lensing enables us to probe distinct stellar populations inside a galaxy.
@@
A370-z6LAE-2 satisfies all of our updated PopIII selection criteria in both the colorācolor diagrams and the SED-based tests, except for the requirement of (Eq.7). The S/N in F410M is only 1.4, which naturally reflects the faintness of the object. As discussed in Sec.Ā 5.1, this specific F410M requirement is designed to exclude contaminants exhibiting a significant Balmer jump, such as the case encountered for GLIMPSE-16043. Although the current position of A370-z6LAE-2 (red open triangle) in FigureĀ 6 lies well away from the region occupied by the exotic Balmer-jump contaminants, the low S/N in F410M permits its inferred location to scatter into that regime. On the other hand, A370-z6LAE-2 is relatively isolated, with no bright neighboring sources, implying that deeper F410M imaging alone, even without the spectroscopy, could robustly establish its color and elevate it to a promising PopĀ III candidate.
@@ Appendix D Survey Area
For each cluster, we compute the source-plane area as a function of magnification , binned logarithmically from to . At each pixel inside the defined region, the magnification is obtained from the lens model (assuming a source redshift of ), and the corresponding source-plane contribution is computed as the image-plane pixel area divided by . Pixels are then accumulated into cumulative magnification bins, such that each bin records the total source-plane area with magnification exceeding a given threshold, . We then convert the relation between the cumulative source-plane area and into a relation between and the effective limiting magnitudes for each field. The limiting depths are obtained by scaling according to the magnification and by adopting the F200W sensitivities reported in the literature for the lensing fields analyzed in this study. For non-lensed general blank fields, we use the detection limits and survey areas compiled in fujimoto2025. The resulting ādepth relations for all fields are presented in FigureĀ 16. These relations are used directly to compute the magnification-corrected survey volume employed in our updated PopĀ III UVLF and CSFRD analysis (SectionĀ 5.3).